# Conditional regression for the Nonlinear Single-Variable Model

### v1

@@ 3.2 Estimation of the distance function and classification accuracy
    Here we assume (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), (LCV), and ($\bm{\omega_{f}}$): by Corollary 8 we are in “thin” slice scenario, i.e. $H_{l,h}>0$, and the distance function simplifies to ${\widehat{\mathrm{dist}}}(x,h)=|\langle x-{\widehat{\mu}_{l,h}^{b}},{\widehat{v}_{l,h}^{b}}\rangle|^{2}+\frac{{\lambda_{d}(\widehat{\Sigma}_{l,h}^{b})}}{{\lambda_{1}(\widehat{\Sigma}_{l,h}^{b})}}\left\|x-{\widehat{\mu}_{l,h}^{b}}\right\|^{2}$. In Algorithm 1, we take the slice ${\widehat{h}_{x}}$, which has the smallest estimated distance, to $x$, as an estimator of the true correct ${h^{\prime}_{x}}$. The following proposition states that the population counterpart ${h_{x}}$ of the nearest index almost equals the correct index ${h^{\prime}_{x}}$.
@@
    Let ${h_{x}}=\operatorname*{{arg\,min}}_{h\in\mathcal{H}_{l}}{\mathrm{dist}}(x,h)$ be the nearest index and define the correct index ${h^{\prime}_{x}}$ be the unique $h\in\mathcal{H}_{l}$ such that $F(x)\in{R_{l,h}}$. Suppose ($\bm{\zeta}$ subG), ($\bm{\omega_{f}}$), and (LCV) hold true. Suppose that $|{R_{l,h}}|\geqslant\max(\sigma_{\zeta},\omega_{f})$ for all $h\in\mathcal{H}_{l}$. Then $|{h_{x}}-{h^{\prime}_{x}}|\leqslant 1$. Moreover, the phenomenon of adjacent misclassification (i.e., $|{h_{x}}-{h^{\prime}_{x}}|=1$) only occurs for points near the boundary of some slices.
@@
    Assume (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), (LCV), and ($\bm{\omega_{f}}$) hold true. If $l\gtrsim\frac{C_{f}\mathrm{len}_{\gamma}}{C_{f}^{\prime}\sigma_{\gamma}}$, and $|{R_{l,h}}|\geqslant\max(\sigma,\omega_{f})$ for all $h\in\mathcal{H}_{l}^{b}$, then the probability of misclassification by at least two slices in part 2.b of Algorithm 1 satisfies
@@ 3.3 Function estimation error corresponding to almost correct classification
    $\widetilde{f}(s)=\mathbb{E}\left[F(X)\ \big{|}\ X\in{S_{l,h}},\ \langle{v_{l,h}^{b}},X\rangle=s\right]$
@@
    $+\underbrace{\left|{\widehat{f}_{j|{v_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)-{\widehat{f}_{j|{\widehat{v}_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)\right|}_{(\Psi)}$
@@
    Consider assumptions (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), (LCV), ($\bm{\omega_{f}}$), and (P). Suppose $f\in\mathcal{C}^{s}$ with $s\in\left[\frac{1}{2},2\right]$. Conditioned on almost correct classification in part 2. b) in Algorithm 1, then we have the following estimates for $n$ is sufficiently large,
    $\operatorname{\mathrm{MSE}}_{(NCA)}\lesssim$ $\left[f\right]_{\mathcal{C}^{{s\wedge 1}}}^{2}(\sigma_{\gamma}\left\|\gamma^{\prime\prime}\right\|C_{f})^{{2(s\wedge 1)}}\max\left(\sigma_{\zeta},\omega_{f},\frac{\mathrm{len}_{\gamma}}{C_{f}^{\prime}l}\right)^{{2(s\wedge 1)}}$
    $\operatorname{\mathrm{MSE}}_{(\Phi)}\lesssim$ $(\left[f\right]_{\mathcal{C}^{{s\wedge 1}}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{{s\wedge 1}}})^{2}(\mathrm{len}_{\gamma}R_{0}^{2}\sigma_{\gamma}^{-2})^{{2(s\wedge 1)}}(d^{2}\log d)^{{s\wedge 1}}\left(\frac{\log n\log l}{nl}\right)^{{2(s\wedge 1)}}$
    $\operatorname{\mathrm{MSE}}_{(B)}\lesssim$ $(\left[f\right]_{\mathcal{C}^{s}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{s}})^{2}C_{f}^{2s}\max\left(\sigma_{\zeta},\omega_{f},\frac{\mathrm{len}_{\gamma}}{C_{f}^{\prime}l}\right)^{2s}j^{-2s}$
    $\operatorname{\mathrm{MSE}}_{(V)}\lesssim$ $\sigma_{\zeta}^{2}\frac{lj\log j}{n}$
@@ 4 Numerical Experiments
    For all figures in this section, we use loglog plots to study how the following quantities decay with the number of samples $n$: the mean squared error $\mathbb{E}[|{\widehat{F}}_{n}(X)-F(X)|^{2}]$, the estimation error of the center along the tangential direction $\mathbb{E}[|\langle{\hat{\mu}}_{l^{*},h}-\mu_{l^{*},h},\gamma^{\prime}_{l^{*},h}\rangle|]$, and the difference between the significant vector and the tangential direction $\mathbb{E}[||{\widehat{v}}_{l^{*},h}-\gamma^{\prime}_{l^{*},h}||]$. On the interval $n\in[10^{5},10^{6}]$, we use least squares linear regression to estimate the decaying rate. We add a dashed vertical line $n=10^{5}$ for those figures to emphasize that the learning rate only corresponds to $n\in[10^{5},10^{6}]$. Because the mean squared error has a term due to the curve approximation error, one should keep in mind that the learning rate for mean squared error in numerical tests here still only indicative of the learning rate for $n\in[10^{5},10^{6}]$ and may be smaller than the min-max optimal exponent $\frac{2s}{2s+1}$ when the mean squared error is dominated by the curve approximation error and the overall mean squared error starts to get saturated, which happens for large $n$. To complicate things further, such curve approximation error depends on curvature, which in turn can depend on the dimension.
@@ 4.1 Example 1: Circular Arcs and verifying Theorem 3
    Theorem 3 claims that the curve approximation error is proportional to $\left\|\gamma^{\prime\prime}\right\|^{{2(s\wedge 1)}}$, which is smaller for curve with smaller curvature: we verify that claim by the following numerical test. We choose $f(t)=\exp(t)$ for $t\in[0,1]$, which has smoothness exponent $s\geq 2$. The external noise follows the normal distribution $\mathcal{N}(0,\sigma_{\zeta}^{2})$ with $\sigma_{\zeta}=0.03$. Figure 4 supports the following conclusions, consistent with Theorem 3 and our analysis: (i) when $\sigma_{\zeta}>0$, then the mean squared error has a nonzero lower bound independent of $n$; (ii) the curve approximation error is proportional to $\left\|\gamma^{\prime\prime}\right\|^{{2(s\wedge 1)}}$; (iii) Corollary 9 states that the decaying rate is $\mathcal{O}(n^{-\frac{1}{2}})$ for $\mathbb{E}[|\langle{\hat{\mu}}_{l^{*},h}-\mu_{l^{*},h},\gamma^{\prime}_{l^{*},h}\rangle|]$ and $\mathbb{E}[||{\hat{\mu}}_{l^{*},h}-\gamma^{\prime}_{l^{*},h}||]$.
@@ 4.2.2 Example: Verifying Theorem 3 and Corollary 9
    In the upper-left plot of Figure 6, we observe that, for fixed $n=10^{5}$, the mean squared error is larger for Meyer helix curves with in ambient dimension $d$. This is because the coefficient before $n^{-\frac{2s}{2s+1}}$ in Theorem 3 is larger for curves with larger $\mathrm{len}_{\gamma}$. In upper right plot, we observe that, by contrast, the mean squared error at $n=2\times 10^{6}$ is smaller for Meyer helix curves in larger ambient dimension $d$. This verifies the statement in Theorem 3 that the curve approximation error is proportional to $\mathrm{reach}_{\gamma}^{-2{s\wedge 1}}$. Some observations about this example: the Meyer helix in smaller $d$ has smaller $\mathrm{len}_{\gamma},\mathrm{reach}_{\gamma}$, and larger curvature, and therefore: (i) the requirement for the number of samples for Theorem 3 is smaller; (ii) the first term in the mean squared error, which is $\mathcal{O}(n^{-\frac{2s}{2s+1}}\log n)$, has a smaller coefficient; (iii) the second term in the bound for the mean squared error in Theorem 3, which is the curve approximation error, has larger magnitude; (iv) if we denote by $n_{1}$ the number of samples such that first term balances the second term in the upper bound for the mean squared error, $n_{1}$ increases with $d$. As a consequence, on the particular interval $n=[10^{5},10^{6}]$ where the decaying rate is calculated, one need to pay attention to the value of $n_{1}$ that determines which term dominates the mean squared error. In particular, the Meyer helix in $d=3$ dimensions has $n_{1}\approxeq 10^{5}$, with the MSE exhibiting good decay for $n\leq n_{1}$, and saturating for $n\geq n_{1}$. In contrast, the Meyer helix in $d=7$ deimsnsions has $n_{1}\approxeq 10^{6}$, and thus we observe a good decay rate on the interval $n\in[10^{5},10^{6}]$. For the Meyer helix in higher dimension, the requirement for the number of samples by Theorem 3 is even larger. There is another phenomenon appearing in this numerical test: we notice that for $d=8,9$, the error increases when $n$ increase from $10^{4}$ to $2\times 10^{4}$ and decreases when $n\geqslant 4\times 10^{4}$. First, this is not a contradiction with Theorem 3 because the requirement for the minimal number of samples is not satisfied for the Meyer helix in dimension $d=8,9$ when $n\leqslant 3\times 10^{4}$. Second, this increase of error for small $n$ is due to the transition from the “wide” slice scenario to the “thin” slice scenario. Further investigation on the average empirical geometric quantity $\frac{1}{l^{*}}\sum_{h=1}^{l^{*}}\widehat{H}_{l^{*},h}$ shows that we have the “wide” slice scenario when $n\leqslant 3\times 10^{3}$ and have the “thin” slice scenario when $n\geqslant 4\times 10^{4}$. For $n\in[3\times 10^{4},4\times 10^{4}]$, the average empirical geometric quantity $\frac{1}{l^{*}}\sum_{h=1}^{l^{*}}\widehat{H}_{l^{*},h}\approx 0$ and thus the slices are roughly isotropic which, as discussed in Section 2.1, prevents an accurate estimate of the significant vector.
@@ 5 Robustness: Performance of Algorithm 1 in general setting
    Suppose (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), and ($\bm{\omega_{f}}$) hold true, together with (SC), with $c_{1},c_{2}$ smaller than a small-enough universal constant. Then, for every $l$ such that $|{R_{l,h}}|\asymp\max(\sigma_{\zeta},\omega_{f}),h\in\mathcal{H}_{l}$, we have
    $\lambda_{1}({\Sigma_{l,h}^{b}})-\lambda_{2}({\Sigma_{l,h}^{b}})\gtrsim C_{f}^{2}|{R_{l,h}}|^{2}$
@@
    Assume that (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), ($\bm{\omega_{f}}$), and (SC) hold true. Fix $l=\frac{C_{Y}R_{0}}{\max(\sigma,\omega_{f})}$. Then, the probability of misclassification by at least two slices in part 2. b) of Algorithm 1 can be bounded by
    $\mathbb{P}\left(\left|{\widehat{h}_{x}}-{h_{x}}\right|\geqslant 2\right)\lesssim ld\exp\left(-c\frac{C_{f}^{6}\max(\sigma_{\zeta},\omega_{f})^{7}n}{C_{Y}R_{0}^{7}d^{3}\sqrt{\log n}}\right)+ln^{-\tau}$
@@ Appendix A Proof of Propositions
    $\mathbb{P}\left(\left\|\widehat{M}\right\|\gtrsim t\right)\lesssim(d_{1}+d_{2})\exp\left(-\frac{cnt^{2}}{Bt+\sigma^{2}}\right)$
@@
    $\mathbb{P}\left(\left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|\gtrsim R_{0}\sqrt{\frac{\delta}{\alpha}}\right)\lesssim d\exp\left(-\frac{c{n_{l,h}^{b}}\delta}{\alpha d}\right),\ \ \text{ where }\delta<\alpha d$
    $\mathbb{P}\left(\left\|{\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}\right\|\gtrsim\beta\frac{R_{0}^{2}}{\alpha}\right)\lesssim d\exp\left(-\frac{c\beta^{2}{n_{l,h}^{b}}}{\alpha^{2}d^{2}}\right),\ \ \text{ where }\beta<\alpha d$
    $\mathbb{P}\left(\left|{\lambda_{1}(\widehat{\Sigma}_{l,h}^{b})}-{\lambda_{1}(\Sigma_{l,h}^{b})}\right|\gtrsim\beta\frac{R_{0}^{2}}{\alpha}\right)\lesssim d\exp\left(-\frac{c\beta^{2}{n_{l,h}^{b}}}{\alpha^{2}d^{2}}\right),\ \ \text{ where }\beta<\alpha d$

### v2

@@ 3.2 Estimation of the distance function and classification accuracy
[…]
@@
    Let ${h_{x}}=\operatorname*{{arg\,min}}_{h\in\mathcal{H}_{l}}{\mathrm{dist}}(x,h)$ be the nearest index and define the correct index ${h^{\prime}_{x}}$ be the unique $h\in\mathcal{H}_{l}$ such that $F(x)\in{R_{l,h}}$. Suppose ($\bm{\zeta}\in\psi_{2}$), ($\bm{\omega_{f}}$), and (LCV) hold true. Suppose that $|{R_{l,h}}|\geqslant\max(\sigma_{\zeta},\omega_{f})$ for all $h\in\mathcal{H}_{l}$. Then $|{h_{x}}-{h^{\prime}_{x}}|\leqslant 1$. Moreover, the phenomenon of adjacent misclassification (i.e., $|{h_{x}}-{h^{\prime}_{x}}|=1$) only occurs for points near the boundary of some slices.
@@
[…]
@@ 3.3 Function estimation error corresponding to almost correct classification
    $\widetilde{f}(s)=\mathbb{E}\left[F(X)\ \big{|}\ X\in{S_{l,h}},\ \langle{v_{l,h}^{b}},X\rangle=s\right]\ \,,$
@@
    $+\underbrace{\left|{\widehat{f}_{j|{v_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)-{\widehat{f}_{j|{\widehat{v}_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)\right|}_{(\Psi)}\ \,.$
@@
[…]
    $\operatorname{\mathrm{MSE}}_{(NCA)}\lesssim$ $\left[f\right]_{\mathcal{C}^{{s\wedge 1}}}^{2}(\sigma_{\gamma}\left\|\gamma^{\prime\prime}\right\|C_{f})^{{2(s\wedge 1)}}\max\left(\sigma_{\zeta},\omega_{f},\frac{\mathrm{len}_{\gamma}}{C_{f}^{\prime}l}\right)^{{2(s\wedge 1)}}\ \,;$
    $\operatorname{\mathrm{MSE}}_{(\Phi)}\lesssim$ $(\left[f\right]_{\mathcal{C}^{{s\wedge 1}}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{{s\wedge 1}}})^{2}(\mathrm{len}_{\gamma}R_{0}^{2}\sigma_{\gamma}^{-2})^{{2(s\wedge 1)}}(d^{2}\log d)^{{s\wedge 1}}\left(\frac{\log n\log l}{nl}\right)^{{2(s\wedge 1)}}\ \,;$
    $\operatorname{\mathrm{MSE}}_{(B)}\lesssim$ $(\left[f\right]_{\mathcal{C}^{s}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{s}})^{2}C_{f}^{2s}\max\left(\sigma_{\zeta},\omega_{f},\frac{\mathrm{len}_{\gamma}}{C_{f}^{\prime}l}\right)^{2s}j^{-2s}\ \,;$
    $\operatorname{\mathrm{MSE}}_{(V)}\lesssim$ $\sigma_{\zeta}^{2}\frac{lj\log j}{n}\ \,;$
@@ 4 Numerical Experiments
    Remark: Throughout this section, we test the performance of Algorithm 1 using theoretically predicted parameters $(l^{*},j^{*})$, and we verify in many aspects that the performance in estimation is consistent with the main theorems. Nevertheless, it is noteworthy that the theoretical argument focuses mainly on the optimal choice $(l^{*},j^{*})$ which depends on many quantities, including parameters of the underlying curve $\gamma$, that are likely to be unknown in real world applications. In practice, the optimal choice for $l,j$ can be found using cross-validation. For example, one can apply the 10-fold cross-validation to the training data to extract the optimal value of parameters $l,j$. In experiments where we used cross-validation we obtained very similar results to those using theoretically optimal parameters.
    For all figures in this section, we use loglog plots to study how the following quantities decay with the number of samples $n$: the mean squared error $\mathbb{E}[|{\widehat{F}}_{n}(X)-F(X)|^{2}]$, the estimation error of the center along the tangential direction $\mathbb{E}[|\langle{\hat{\mu}}_{l^{*},h}-\mu_{l^{*},h},\gamma^{\prime}_{l^{*},h}\rangle|]$, and the difference between the significant vector and the tangential direction $\mathbb{E}[||{\widehat{v}}_{l^{*},h}-\gamma^{\prime}_{l^{*},h}||]$. On the interval $n\in[10^{5},10^{6}]$, we use the least squares linear regression to estimate the decaying rates; note that this estimation is not ideal both in the range of small $n$, where our theory may not apply, and for large $n$, due to the error saturation predicted in the additive term in the main theorem. We add a dashed vertical line $n=10^{5}$ for those figures to emphasize that the learning rate only corresponds to $n\in[10^{5},10^{6}]$. Because the mean squared error has a term due to the curve approximation error, one should keep in mind that the learning rate for mean squared error in numerical tests here still only indicative of the learning rate for $n\in[10^{5},10^{6}]$ and may be smaller than the min-max optimal exponent $\frac{2s}{2s+1}$ when the mean squared error is dominated by the curve approximation error and the overall mean squared error starts to get saturated, which happens for large $n$. To complicate things further, such curve approximation error depends on curvature, which in turn can depend on the dimension.
@@ 4.1 Example 1: Circular Arcs and verifying Theorem 3
    Theorem 3 claims that the curve approximation error is proportional to $\left\|\gamma^{\prime\prime}\right\|^{{2(s\wedge 1)}}$, which is smaller for curve with smaller curvature: we verify that claim by the following numerical test. We choose $f(t)=\exp(t)$ for $t\in[0,1]$, which has smoothness exponent $s\geq 2$. The external noise follows the normal distribution $\mathcal{N}(0,\sigma_{\zeta}^{2})$ with $\sigma_{\zeta}=0.03$. Figure 4 supports the following conclusions, consistent with Theorem 3 and our analysis: (i) When we have nontrivial external noise $\sigma_{\zeta}>0$, then the mean squared error has a nonzero lower bound independent of $n$; Indeed, in the upper-left plot, we observe that as $n$ increases, the MSE decays at a slower rate and barely changes when $n\approx 10^{6}$; This verifies the existence of constant approximation error. Note that the average decaying rate for $n\in[10^{5},10^{6}]$ is estimated using least squares. (ii) The curve approximation error is proportional to $\left\|\gamma^{\prime\prime}\right\|^{{2(s\wedge 1)}}$; In fact, in the left plot, we observe that the value of MSE at $n=10^{6}$ is larger for curves with higher curvature and smaller for curves for smaller curvature. In the right plot, we further study their relation between the value of MSE when $n=10^{6}$ and the magnitude of curvature. We also use the least squares to estimate the growth rate and the estimated growth rate is consistent with the theoretical counterpart. (iii) Corollary 9 states that the decaying rate is $\mathcal{O}(n^{-\frac{1}{2}})$ for $\mathbb{E}[|\langle{\hat{\mu}}_{l^{*},h}-\mu_{l^{*},h},\gamma^{\prime}_{l^{*},h}\rangle|]$ and $\mathbb{E}[||{\hat{\mu}}_{l^{*},h}-\gamma^{\prime}_{l^{*},h}||]$.
@@ 4.2.2 Example: Verifying Theorem 3 and Corollary 9
    In the left plot of Figure 6, we observe that, for fixed $n=10^{5}$, the mean squared error is larger for Meyer helix curves with increasing ambient dimension $d$. This is because the coefficient before $n^{-\frac{2s}{2s+1}}$ in Theorem 3 is larger for curves with larger $\mathrm{len}_{\gamma}$. In the right plot, we observe that, in contrast, the mean squared error at $n=2\times 10^{6}$ is smaller for Meyer helix curves in larger ambient dimension $d$. This verifies the statement in Theorem 3 that the curve approximation error is proportional to $\mathrm{reach}_{\gamma}^{-2{s\wedge 1}}$. Some observations about this example: the Meyer helix in smaller $d$ has smaller $\mathrm{len}_{\gamma},\mathrm{reach}_{\gamma}$, and larger curvature, and therefore: (i) the requirement for the number of samples for Theorem 3 is smaller; (ii) the first term in the mean squared error, which is $\mathcal{O}(n^{-\frac{2s}{2s+1}}\log n)$, has a smaller coefficient; (iii) the second term in the bound for the mean squared error in Theorem 3, which is the curve approximation error, has larger magnitude; (iv) if we denote by $n_{1}$ the number of samples such that first term balances the second term in the upper bound for the mean squared error, $n_{1}$ increases with $d$. As a consequence, on the particular interval $n=[10^{5},10^{6}]$ where the decaying rate is calculated, one need to pay attention to the value of $n_{1}$ that determines which term dominates the mean squared error. In particular, the Meyer helix in $d=3$ dimensions has $n_{1}\approxeq 10^{5}$, with the MSE exhibiting good decay for $n\leq n_{1}$, and saturating for $n\geq n_{1}$. In contrast, the Meyer helix in $d=7$ deimsnsions has $n_{1}\approxeq 10^{6}$, and thus we observe a good decay rate on the interval $n\in[10^{5},10^{6}]$. For the Meyer helix in higher dimension, the requirement for the number of samples by Theorem 3 is even larger. There is another phenomenon appearing in this numerical test: we notice that for $d=8,9$, the error increases when $n$ increase from $10^{4}$ to $2\times 10^{4}$ and decreases when $n\geqslant 4\times 10^{4}$. First, this is not a contradiction with Theorem 3 because the requirement for the minimal number of samples is not satisfied for the Meyer helix in dimension $d=8,9$ when $n\leqslant 3\times 10^{4}$. Second, this increase of error for small $n$ is due to the transition from the “wide” slice scenario to the “thin” slice scenario. Further investigation on the average empirical geometric quantity $\frac{1}{l^{*}}\sum_{h=1}^{l^{*}}\widehat{H}_{l^{*},h}$ shows that we have the “wide” slice scenario when $n\leqslant 3\times 10^{3}$ and have the “thin” slice scenario when $n\geqslant 4\times 10^{4}$. For $n\in[3\times 10^{4},4\times 10^{4}]$, the average empirical geometric quantity $\frac{1}{l^{*}}\sum_{h=1}^{l^{*}}\widehat{H}_{l^{*},h}\approx 0$ and thus the slices are roughly isotropic which, as discussed in Section 2.1, prevents an accurate estimate of the significant vector.
@@ 5 Robustness: Performance of Algorithm 1 in general setting
[…]
    $\lambda_{1}({\Sigma_{l,h}^{b}})-\lambda_{2}({\Sigma_{l,h}^{b}})\gtrsim C_{f}^{2}|{R_{l,h}}|^{2}\ \,.$
@@
[…]
    $\mathbb{P}\left(\left|{\widehat{h}_{x}}-{h_{x}}\right|\geqslant 2\right)\lesssim ld\exp\left(-c\frac{C_{f}^{6}\max(\sigma_{\zeta},\omega_{f})^{7}n}{C_{Y}R_{0}^{7}d^{3}\sqrt{\log n}}\right)+ln^{-\tau}\ \,.$
@@ Appendix A Proof of Propositions
    $\mathbb{P}\left(\left\|\widehat{M}\right\|\gtrsim t\right)\lesssim(d_{1}+d_{2})\exp\left(-\frac{cnt^{2}}{Bt+\sigma^{2}}\right)\ \,.$
@@
    $\mathbb{P}\left(\left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|\gtrsim R_{0}\sqrt{\frac{\delta}{\alpha}}\right)\lesssim d\exp\left(-\frac{c{n_{l,h}^{b}}\delta}{\alpha d}\right),\ \ \text{ where }\delta<\alpha d\ \,;$
    $\mathbb{P}\left(\left\|{\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}\right\|\gtrsim\beta\frac{R_{0}^{2}}{\alpha}\right)\lesssim d\exp\left(-\frac{c\beta^{2}{n_{l,h}^{b}}}{\alpha^{2}d^{2}}\right),\ \ \text{ where }\beta<\alpha d\ \,;$
    $\mathbb{P}\left(\left|{\lambda_{1}(\widehat{\Sigma}_{l,h}^{b})}-{\lambda_{1}(\Sigma_{l,h}^{b})}\right|\gtrsim\beta\frac{R_{0}^{2}}{\alpha}\right)\lesssim d\exp\left(-\frac{c\beta^{2}{n_{l,h}^{b}}}{\alpha^{2}d^{2}}\right),\ \ \text{ where }\beta<\alpha d\ \,.$